If y is directly proportional to x and y=8 when x=10 what is the value of ywhen x=5

Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
y = Kx
8 = K(10)
K = 8/10
---
y = (8/10)x
x = 5
y = (8/10)5
y = 4
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

If y is directly proportional to x
y = kx
-----
Solve for "k" using "y=8 when x=10"
8 = k*10
k = 4/5
----
Equation:
y = (4/5)x
----
What is the value of y when x=5?
---
y = (4/5)(5)
--
y = 4
==================
Cheers,
Stan H.
------------------